OpenStudy (anonymous):

How are you going to integrate f(x) = integral of -(sin(y)) dy/ (3tany-2cosy) ?

OpenStudy (anonymous):

i simplified it into integral of (-sinycosy)dy / (2sin^y +3 siny-2)

OpenStudy (dumbcow):

Not sure on the steps..but here is the solution its pretty long http://www.numberempire.com/integralcalculator.php?function=%28-sin%28x%29%29%2F%283*tan%28x%29-2*cos%28x%29%29&var=x&answers=&__utma=164170016.165112442.1306893325.1312516703.1312708626.8&__utmz=164170016.1312708628.8.7.utmcsr%3Dgoogle |utmccn%3D%28organic%29|utmcmd%3Dorganic|utmctr%3Dintegral+calculator&screen_height=768&__utmb=164170016.1.10.1312708628&__utmc=164170016

OpenStudy (anonymous):

Wolfram has some good steps for the integral, I am just wondering if it is possible to get a sensible trig expression from the original...

OpenStudy (anonymous):

I think we can....

OpenStudy (dumbcow):

i can simplify it to this...then maybe partial fractions ? \[-\frac{\sin x *\cos x}{(2\sin x +1)(\sin x -2)}\]

OpenStudy (anonymous):

Think I something better, maybe u check me, divide original top and bottom by 2 siny.

OpenStudy (dumbcow):

ok

OpenStudy (anonymous):

Might have messed it up (not awake yet:-) but -2Tan^2y/(3Tan^2y-2)

OpenStudy (dumbcow):

wait how do you get tan^2 ? divide top by 2sin --> -1/2 divide bottom by 2sin --> (3/2cos) - cot

OpenStudy (anonymous):

Ooops, told u not awake:-) Still, I thing something is there...

OpenStudy (anonymous):

Think..

OpenStudy (dumbcow):

ok put it together... \[\frac{\tan x}{3\sin x - 2}\]

OpenStudy (anonymous):

That is already easier, right..?

OpenStudy (dumbcow):

i dunno, what can we substitute here ??

OpenStudy (anonymous):

Maybe something /u^2+1...

myininaya (myininaya):

i did not see to do this

myininaya (myininaya):

wolfram will show you the steps

myininaya (myininaya):

they do u=tan(y/2) and then use partial fractions

OpenStudy (anonymous):

It is the partial fractions I am trying to get out of..... I'm sure there is a way, just haven't seen it yet...:-)

OpenStudy (anonymous):

Maybe try to use sec instead...

OpenStudy (anonymous):

eewww calculus.

OpenStudy (anonymous):

Well, trig at the moment....

myininaya (myininaya):

actually this problem does look ewwwie

OpenStudy (anonymous):

ewwwie - another one for me to learn..:-)?

OpenStudy (anonymous):

urban dictionary says it is an ex of some kind...

OpenStudy (dumbcow):

ok i just worked it out using the simplified expression i posted above...partial fractions does work u = sin(x) - 2 du = cos(x)

OpenStudy (anonymous):

There u go, I don't like partial fractions though, evil...

OpenStudy (dumbcow):

haha :)

OpenStudy (dumbcow):

my steps aren't as messy as wolframs...lol

OpenStudy (dumbcow):

i don't like using sec and tan if i can help it, just stick with sin and cos

OpenStudy (anonymous):

manoranjan just told me in chat that eewwie is dental floss..

OpenStudy (dumbcow):

yeah i'm sure there are a lot of things that can be defined as "eewwie",

OpenStudy (anonymous):

Or ewwwie..

OpenStudy (dumbcow):

so wheres "thebeefbowl" ? nice name btw hopefully all this will be of some help, i don;t think i want to write out my whole solution

OpenStudy (anonymous):

Sure he can do with this info...

OpenStudy (anonymous):

Ok \[\frac{-siny}{3tany -2cosy}dy\] \[Put\:\:\:\:\:\:\: t = siny \]\[\frac {dt}{cosy} = dy\] \[\frac{-t}{3t -2 + 2t^2}dt\] Now, \[ \lambda \frac{d(3t -2 + 2t^2)}{dt} + \mu = -t\]

OpenStudy (anonymous):

Ooooh....!

OpenStudy (anonymous):

Course, the trig would be easy if it wasn't for the 3 and the 2....

OpenStudy (anonymous):

Using this we Have \[3\lambda + 4t\lambda +\mu = -t\]Compare coefficients \[3\lambda + \mu = 0\] \[4\lambda = -1\]

OpenStudy (anonymous):

G, epsilons next...:-)

OpenStudy (anonymous):

Now when we Integrate it \[f(t) = 3t - 2 +2t^2\] \[\frac{\lambda f'(t) + \mu }{f(t)} dt\] lol@estudier I am just trying to do it thats all

OpenStudy (anonymous):

No, no...I like it, keep going...:-)

OpenStudy (anonymous):

\[\lambda \frac{f'(t)}{f(t)} dt + \frac{\mu}{f(t)}dt\] Now first part integral is \[\lambda log(f(t))\]

OpenStudy (anonymous):

I keep getting kicked off...

OpenStudy (anonymous):

Now all we have to do is the second part \[\mu \frac{1}{3t - 2 + 2t^2} dt \]

OpenStudy (anonymous):

\[3t - 2 + 2t^2 = ((t + 2) (2t -1)\] \[ \frac{1}{3t - 2 +2t^2} = \frac{A(2t -1) +B(t +2)}{3t - 2 + 2t^2 } \]

OpenStudy (anonymous):

Now, Compare coefficients The Answer should be \[\lambda log(3t + 2t^2 -2 ) + \mu A log(t+2) + \mu B log(2t -1)\]

OpenStudy (anonymous):

ah I am not putting coefficients or t ...you should do algebra on you own

OpenStudy (anonymous):

I think I did wrong ...wolfram is showing something very complex

OpenStudy (anonymous):

Don't pay attention to Wolfram on this one....

OpenStudy (anonymous):

Maybe your right

OpenStudy (anonymous):

-sin2y /2 (siny+2)(2siny-1)....Best I can get, I think.

OpenStudy (anonymous):

That's from multiplying original integral by cosy top and bottom and u can factor the bottom as a quadratic in sin y...

OpenStudy (anonymous):

Mind u I just noticed in the originall question he wants to integrate the integral,lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

haha thanks everyone. i got the idea now. and my question is wrong. it should be the integral o f the function only not the integral of integral :D

OpenStudy (anonymous):

my answer is 1/10 ln|2sinx -1| + 2/5 ln |sinx+2| +C